If r is the root of f(x)=x^2-4 and t is the root of g(x)= x^2 -25 what is the value of lr-til
Answer by stanbon(75887)

If r is the root of f(x)=x^2-4 and t is the root of g(x)= x^2 -25 what is the value of lr-til
-----------
r= 2 or -2
t=5 or -5
|r-ti| = |2-5i| = sqrt(2^2+5^2)= sqrt(29)
The other combinations of values of r and t give the same result.
Cheers,
Stan H.